# Selection of Filters for Photonic Crystal Spectrometer Using Domain-Aware Evolutionary Algorithms

### v1

@@ Abstract
This work addresses the critical challenge of optimal filter selection for a novel trace gas measurement device. This device uses photonic crystal filters to retrieve trace gas concentrations prone to photon and read noise. The filter selection directly influences accuracy and precision of the gas retrieval and therefore is a crucial performance driver. We formulate the problem as a stochastic combinatorial optimization problem and develop a simulator mimicking gas retrieval with noise. The objective function for selecting filters reducing retrieval error is minimized by the employed metaheuristics, that represent various families of optimizers. We aim to improve the found top-performing algorithms using our novel distance-driven extensions, that employ metrics on the space of filter selections. This leads to a novel adaptation of the UMDA algorithm, we call UMDA-U-PLS-Dist, equipped with one of the proposed distance metrics as the most efficient and robust solver among the considered ones. Analysis of filter sets produced by this method reveals that filters with relatively smooth transmission profiles but containing high contrast improve the device performance. Moreover, the top-performing obtained solution shows significant improvement compared to the baseline.
keywords:
Photonic crystal, Trace Gas Measurements, Spectrometer, Metaheuristic, Evolutionary Computation, Domain-Aware Optimization, Stochastic Combinatorial Optimization, Simulation Optimization
@@ 1 Introduction
Figure 1: Graphical abstract. Panel 1 illustrates the satellite instrument with a photonic crystal spectrometer, where the instrument model contains photonic crystal transmission filters. Panel 2 shows an example of selecting a diverse set of high-quality solutions. Panel 3 depicts how the optimized selection of the filters outperforms the baseline.
@@
As global climate change severely impacts our world, there is an increasing demand to monitor trace gases with a high spatial resolution and accuracy. At the same time, these instruments and satellites need to be compact in order to have constellations for short revisit times [1]. In this study we explore a new spectrometer instrument concept to detect trace gases where photonic crystal filters directly placed on a detector replace traditional diffraction-based optical elements [2]. We give at first a high-level description of the main challenge in the implementation of this concept, while providing more details about the spectrometer in Section 2. Trace gasses in the atmosphere slightly change the spectrum of the transmitted light due to absorption at specific wavelengths. The instrument is able to perform trace gas retrieval by measuring the total amount of reflected light each filter transmits for a certain ground pixel. By analyzing the measured intensities, it is possible determine the concentration of trace gases. The performance of the instrument, measured as accuracy and precision in gas retrieval, depends on the signal-to-noise ratio and significantly on the selected filters. Therefore, a key challenge is to select the filters that will most accurately measure the trace gas concentration. To achieve this, we simulated how 2D photonic crystals respond to light in various shapes and sizes, creating a library of about 5000 different filters. Given that the considered instrument can have at maximum 640 filters, we reduce the problem to the selection of an optimal combination of filters from the created library. In this study, we are specifically focusing on detecting methane as a test case. Therefore, we can measure the instrument’s performance by comparing the detected methane concentration with the actual concentration.
@@
This paper addresses the described problem of selecting from a large library of approximately $L\approx 5\cdot 10^{3}$ filters a combination that optimizes the instrument’s performance in methane retrieval. Unlike a set, where each element is unique, filters in this selection can be chosen multiple times, i.e., a solution candidate is a multiset with $N$ elements chosen from $L$. For the remainder of this paper, we refer to this core problem as Optimal Filter Selection (OFS). The problem is addressed by an optimization algorithm that tackles both challenges: a great number of possible solutions and noisy data. The latter is a special challenge, as many evaluations of different filter sets can result in outliers, eroding the convergence of the optimization algorithm. Therefore, it is most natural to formulate OFS as a stochastic combinatorial optimization problem.
@@ 1.1 Combinatorial Optimization
The field of combinatorial optimization [3, 4] studies algorithms that aim to find an object (or a set of objects) that minimize a defined cost criterion. Given that these algorithms were created to solve specific tasks, we begin by presenting an overview of common combinatorial problems. We emphasize that big instances of those tasks demand excruciatingly big computational resources by providing their time complexities. One of the most common examples of a combinatorial optimization problem is the Traveling Salesperson Problem (TSP) [5, 6], where the desired solution is a path in a weighted graph, and the cost criterion to be minimized is the sum of the weights of all the edges on the path. TSP is $\mathcal{NP}$-hard [7] on par with most of the practically relevant combinatorial optimization problems (COPs). Other classical examples of COPs include the weakly $\mathcal{NP}$-hard knapsack problem [8], which is arguably one of the simplest $\mathcal{NP}$ problems [9]. However, solving this problem’s particular instances still takes significant time for numerical solvers considered in [10]. Moreover, many generalizations and variations of this problem, inspired by real-world applications, are considered in the literature [11].
The max cut [12] problem is an example of an $\mathcal{NP}$-complete problem that is sometimes considered one of the most difficult in combinatorial optimization [13]. Another example of COP is a very practical portfolio optimization [14], which is highly relevant in the financial investment domain. An interested reader can find other classical instances of COPs in [4], where a high-level classification of the algorithms that address such problems is proposed.
According to this classification, each algorithm falls into one of three categories: 1) computation of the exact optimal solution, mostly studied in [3, 4]; 2) Computation of approximate solutions [15]; 3) State-of-the-art numerical solvers, reviewed in [16, 17]. Algorithms in the former two categories are conventional combinatorial optimization methods (CCOMs). They search via mathematical or dynamic programming and employ heuristic ideas to disregard intuitively sub-optimal solutions. Particular instances of CCOMs are implemented to solve a specific known COP and therefore heavily depend on the known structure of the addressed problem.
The third category of algorithms includes methods that apply nature-inspired meta-heuristics [18] and machine learning [19] approaches to find solution candidates with sufficiently small costs. While these optimizers do not guarantee to locate the optimal solution, they are less dependent on the constraint nature of the problem and so can be applied to different classes of COPs without major modifications. Those and other relevant research directions are systematically studied in this recent survey [20].
@@ 1.2 Stochastic Combinatorial Optimization
Stochastic combinatorial optimization problems (SCOPs) introduce an additional layer of complexity: the cost associated with a specific solution candidate exhibits randomness, meaning it may vary upon reevaluation. This property is especially relevant for engineering applications as the cost is estimated by a simulation of a real-world process, for example, see Part IV in [21]. Application of an algorithm to a SCOP that relies on a simulation is called simulation optimization (SO) in the literature [21, 22, 23], since a closed-form mathematical expression is not available to describe the optimized function. SO can be classified as a particular case of a more general Black-Box Optimization task [24]. CCOMs can be applied as simulation optimizers, for example, see Chapter 2 of [22]. However, rapidly developing areas of relevant meta-heuristic and machine learning methods [25] contain alternative promising algorithms. As shown in Table 1 of [25], choosing an optimizer in the latter category was advantageous in a number of practical use cases. See the survey [26] for more examples and a fresh systematic review of the related publications.
@@ 1.3 Addressing OFS problem
While CCOMs for OFS might exist, the complexity of deriving them may not be practical. Therefore, we leverage the efficiency of numerical methods, such as meta-heuristics and machine learning, to obtain an approximate solution to OFS. Given the vast number of existing general-purpose numerical methods, we use the latest (known to us) comprehensive survey [23]. This survey provides a clear description of various methods applicable to SO, helping us in selecting optimizers for our OFS task. Independently of our choice, it is guaranteed by the No Free Lunch (NFL) theorem that none of the selected optimizers is the best on average for a big enough class of problems, meaning, a class of problems closed under the permutations of the search space [27]. Therefore, knowing the best optimizer for OFS in advance is intractable. Hence, we follow the approach suggested in Section 2.7 of [17]. We conduct experiments where we repeatedly approximate the solution of the OFS problem using every chosen algorithm. This allows us to statistically rank the performance of each algorithm and identify the most suitable ones for our specific problem. We call the top-performing solvers the leading algorithms.
@@
    We reformulate OFS as a tractable instance of SCOP, compare meta-heuristics that represent various classes of SCOP solvers, and select the solvers with the best performance;
@@
    We solve the following non-trivial technical problems that emerge in stage 2: propose a heuristic algorithm for solving the inverse of the Linear Assignment Problem, required to develop the mutation operator; propose a robust method which we call which we call UMDA-U-PLS-Dist to guide the search leveraging a given distance metric;
@@
    We study the obtained high-performing solutions to OFS problem.
@@ 2 Description of the Trace Gas Measurement Device
The Trace Gas Measurement Device (TGMD) consists of an imaging telescope, where photonic crystal filters, acting as transmission filters, are placed directly on the detector. As the instrument flies over the earth, the spectrum of light of a single ground pixel is projected on to pixels with different optical filters. Each pixel then measures the total transmitted intensity for this filter. Effectively, the spectrum of light is mapped to the (compressed) bases of the transmission profiles of the filters. The trace gas concentration can then be retrieved directly from the different measured integrated intensities. In this study we focus on methane retrieval as an initial test case. The detector in question has 512$\times$640 pixels and is orientated such that the 512 pixels are perpendicular to the flight direction and the 640 pixels are along the flight direction. A single ground resolution element is then observed by 640 pixels as the instrument flies over the earth. Therefore, there can be at maximum 640 filters placed on the detector, one for each pixel. However, current practical limitations such as manufacturing constraints on the distance between the filters and the detector and electromagnetic interactions between neighboring filters limits the number to 16 different filters. This number of filters still contains redundancy in case of in-flight variations of the filter response and unexpected biases due to mismatches in the atmosphere model and imperfect supporting (weather) data, such as wind speeds.
As the light is incident on the filters, different integrated intensities for each filter are measured for a single ground resolution element. From this detector data, the gas concentrations can be retrieved. The exact retrieval algorithm, based on a linearized radiative transfer atmosphere model [28], is out of scope for this work. In short, the trace concentration is used as input for the atmosphere model, see Figures 2 and 3. The atmosphere model produces the spectrum reflected by the earth which is used as input for the instrument model, taking into account the specific filters, which computes the measured values to which the photon and read noise is added. From these noisy measurement values, an inverse model of the instrument and atmosphere are used to retrieve the trace gas concentration using a fitting routine. By performing this fit for different noisy realizations, the fit precision of the trace gas concentration is assessed. This precision is the performance indicator of the instrument concept and should be as low as possible.
In general, the achieved gas concentration retrieval precision depends on many factors such as optical aperture and ground resolution. The specifications and parameters defining the final performance of the instrument are chosen to be reasonable and achievable. Therefore the achieved precision after filter selection optimization is a representative and likely achievable retrieval precision. Currently, the expected retrieval error of TANGO (Twin Anthropogenic Greenhouse gas Observers), a ESA Scout mission to measure methane which consists of a traditional state-of-the-art methane retrieval instrument, is below 1.0% [29]. However, the optical design of the proposed instrument using the photonic crystals is expected to be 5-10 times smaller.
Prior to the development of optimization algorithms we constructed a reasonable filter selection, which we will refer to as SRON baseline. As this initial solution we choose filters based on the sharpness of the spectral transmission features. We used the second moment of the Fourier transform of the transmission profile to rank all filters. Filters with sharp spectral features would result in large amplitudes for the high frequency components. From this ranked list 160 filters are chosen with the largest second moment and this filter set is used as a reference solution. As this method is not likely to include good filters, the solution includes more than 16 filters.
@@ 3.1 Search Space
The search space is defined as sequences of size $N$ of the filters from the library with possible repetitions of the same filter. Notably, the permutation of the filters in one sequence does not change the selection and so does not influence the value of the objective function. This detail will be taken into account later in Section 5.1. Let us denote this space as $\mathscr{F}$. We assume that the cost criterion is given and define the optimization problem, which depends on this criterion.
We start with an intuitive representation of solutions commonly used in Portfolio Optimization [14]. Each filter from the library is taken a certain (possibly zero) number of times so we can represent a candidate solution by a vector $\bm{w}=(w_{1},w_{2},\dots,w_{L})$ of size $L$. The component $w_{i}$ in this vector means that the filter with the number $i$ is taken $w_{i}$ times. All such vectors belong to the $L$-dimensional integer space, which we denote as $\mathscr{F}=\mathbb{N}_{0}^{L}$, where $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup{0}$.
Assume that we are given an objective function $\mathfrak{f}:\mathscr{F}\to\mathbb{R}$, which assesses the quality of the considered sequence of filters. Let us consider all filters in the library as a vector $\bm{l}=\left(l_{1},l_{2},\dots,l_{L}\right)$ of length $L$. If $n$ is an integer and $1\leq i\leq L$, then notation $n\cdot l_{i}$ means that filter number $i$ is included $n$ times to the solution of OFS. The vector $\bm{w}=\left(w_{1},w_{2},\dots,w_{L}\right)$ defines that the filter $i$ is included $w_{i}$ times to the solution of OFS. We look for the weight vector $\bm{w}\in\mathbb{N}_{0}^{L}$, so it is convenient to define a function $\mathfrak{F}:\mathbb{N}_{0}^{L}\to\mathbb{R}$ such that $\mathfrak{F}(\bm{w})=\mathfrak{f}\left(w_{1}\cdot l_{1},w_{2}\cdot l_{2},\ldots,w_{L}\cdot l_{L}\right)$.
Problem Formulation 1 (Optimal Filter Selection (OFS) problem in portfolio optimization form).
Find a vector $\bm{w}^{*}\in\mathbb{N}_{0}^{L}$ which satisfies all the conditions of Eq. (1).
$\mathfrak{F}(\bm{w}^{*})=\min\limits_{\bm{w}\in\mathbb{N}_{0}^{L}}{\mathfrak{F}\!\left(\bm{w}\right)}$ (1a)
$N=\sum\limits_{i=1}^{L}w_{i}$ (1b)
$M=|\left\{w_{i}\neq 0\,\middle|\,i=1,2,\dots,L\right\}|$ (1c)
Eq. (1a) defines the optimization problem, eq. (1b) demands the sum of the weights to be exactly $N$, meaning $N$ filters are used in total, and eq. (1c) constraints the number of different filters in the selection to be $M=16$, as explained in Section 2.
The optimization problem defined in Eq. (1) constitutes a significant challenge for state-of-the-art combinatorial optimization algorithms due to its inherent complexity. This is due to the combination of high dimensionality and the presence of equality constraints, as seen in Eq. (1b). Now we reformulate this statement in order to simplify the optimization problem into a fully equivalent problem.
Let us note that these constraints lead to sparse solutions – vectors with many zero entries. Indeed, all vectors $\bm{w}$ that satisfy constraints Eq. (1b) have at least $L-N$ zeros. Since $L\approx 10\cdot N$, such vectors $\bm{w}$ are very sparse. Let us represent a sequence of selected filters as a vector $\bm{x}=(x_{1},x_{2},\dots,x_{N})\in\{1,\ldots,L\}^{N}$, where every $x_{i}\in\{1,\ldots,L\}$ denotes a particular filter. Such a representation eliminates zero components and reduces the dimensionality of the search space, therefore simplifying the problem. For every such vector $\bm{x}$ there exists a unique vector $\bm{w}\in\mathbb{N}_{0}^{L}$ that satisfies constraints Eq. (1b), since $w_{i}=|\left\{x_{j}=i\;\,\middle|\,\;x_{j}\in\bm{x}\right\}|$. And vice versa, $\bm{x}=(x_{1},x_{2},\dots,x_{N})=\left(w_{i_{1}}\cdot x_{1},w_{i_{2}}\cdot x_{2},\dots,w_{i_{k}}\cdot x_{k}\right)$ such that $i_{1},i_{2},\dots,i_{k}$ are the only non-zero components of vector $\bm{w}$. This mapping $\Phi_{1}:\{1,\ldots,L\}^{N}\rightarrow\mathbb{N}_{0}^{L}$, $\Phi_{1}(\bm{x})=\bm{w}$ is depicted in Figure 4 for convenience of reading. We will use $\Phi_{1}$ to map the vector $\bm{x}$ with reduced dimensionality to the original space, where function $\mathfrak{F}$ is defined.
[…]
Figure 4: Visualization of the mapping $\Phi_{1}:\{1,\ldots,L\}^{N}\rightarrow\mathbb{N}_{0}^{L}$, $\Phi_{1}(\bm{x})=\bm{w}$. The part colored in red depicts the application of $\Phi_{1}$ to a vector encoded in the space $\left\{1,\dots,L\right\}^{N}$. The blue part represents the backward mapping from the space $\mathbb{N}_{0}^{L}$.
Since it is allowed to have the same filter multiple times, vectors $\bm{x}$ may have repeating components. The new dimensionality $N$ can be reduced further if we encode several repeating filters $i$ by one filter $i$. However, we must keep track of the number of times the filter was used in the vector $\bm{x}$. In order to make it, we formulate the following Assumption 1.
Assumption 1.
Let an integer $M$ be the number of different filters, constant $k\coloneqq\lfloor N/M\rfloor$ and constant $r\coloneqq N-k\cdot M$. Then there exists a vector
$(\displaystyle\bm{x}^{*}={}($$(k+1)\cdot x_{1}^{*},(k+1)\cdot x_{2}^{*},\dots,(k+1)\cdot x_{r}^{*},$
$k\cdot x_{r+1}^{*},k\cdot x_{r+2}^{*},\dots,k\cdot x_{M}^{*})$
such that $\mathfrak{F}\circ\Phi_{1}(\bm{x}^{*})$ is sufficiently close to $\min\left\{\mathfrak{F}\circ\Phi_{1}(\bm{x})\,\middle|\,M=\left|\left\{x_{1},x_{2},\dots,x_{N}\right\}\right|\right\}$.
Assumption 1 argues that we can repeat filters almost evenly and do not drastically lose the quality of the solution. This allows us to reduce dimensionality from $N$ to $M$ filters for the given in advance $M$. Given such vector $\bm{x}=(x_{1},x_{2},\dots,x_{M})$ of filters, we can obtain an $N$ dimensional vector by repeating every filter multiple times, as shown in Assumption 1. Let us denote the mapping that performs this transformation as $\Phi_{2}$, which is visualized in Figure 5. Note that $\Phi_{2}$ is not a bijection. In this work, we take $M=16$ because it is the smallest constant such that the obtained solutions can satisfy the constraint in Eq. (1c). Limiting the number of filter types provides a practical benefit: reduced manufacturing costs due to a less complex production process for a smaller variety of different filters, as discussed in Section 2.
[…]
Figure 5: Visualization of the mapping $\Phi_{2}:\{1,\ldots,L\}^{M}\rightarrow\{1,\ldots,L\}^{N}$. The red part of the solution represents the portion where the same filter is duplicated $k+1$ times, blue portion corresponds to $k$ repetitions. Constants are defined as follows: $k=\lfloor N/M\rfloor,r=N-k\cdot M.$
Finally, we are ready to define the OFS problem in the form addressed in this work. For convenience, we will encode every filter with an integer from 0 to $L-1$ and say that $\mathbb{L}=\mathbb{N}_{0}$.
Problem Formulation 2 (OFS problem in the form addressed in this work).
@@
For the given constant $M$, let =$\mathscr{L}=\left\{\bm{x}=(x_{1},x_{2},\dots,x_{M})\,\middle|\,x_{i}\in\mathbb{L}\right\}$ be the space of candidate solutions. For function $\mathrm{F}:\mathscr{L}\to\mathbb{R}$ such that $\mathrm{F}(\bm{x})=\mathfrak{F}\circ\Phi_{1}\circ\Phi_{2}(\bm{x})$ [^1] Notation $f\circ g(x)$ stands for $f(g(x))$., OFS problem is to find $\bm{x}^{*}\in\mathscr{L}$ which satisfies the Eq. (2).
@@
The OFS in the form defined in Formulation 1 is convenient for integer programming, since every component in a candidate solution can be considered independently from others. From the practical perspective, the constraint Eq. (2b) is soft, meaning it can be insignificantly violated. The cardinality of the space $\mathscr{L}$ equals $L^{M}$, which is large in our particular instance of OFS: $L\approx 5000,M=16,L^{M}\approx 10^{59}$. However, all permutations of components in a candidate solution $\bm{x}\in\mathscr{L}$ result in the same vector $\bm{w}=\Phi_{1}\circ\Phi_{2}(\bm{x})$, so they all have the same associated cost. We show how we can take advantage of this property when doing numerical optimization in Section 5.

### v2

@@ Abstract
This work addresses the critical challenge of optimal filter selection for a novel trace gas measurement device. This device uses photonic crystal filters to retrieve trace gas concentrations prone to photon and read noise. The filter selection directly influences the accuracy and precision of the gas retrieval and therefore is a crucial performance driver. We formulate the problem as a stochastic combinatorial optimization problem and develop a simulator modeling gas retrieval with noise. Metaheuristics representing various families of optimizers are used to minimize the retrieval error objective function. We improve the found top-performing algorithms using our novel distance-driven extensions, which employ metrics on the space of filter selections. This leads to a novel adaptation of the Univariate Marginal Distribution Algorithm (UMDA), which we call the Univariate Marginal Distribution Algorithm Unified by Probabilistic Logic Sampling driven by Distance (UMDA-U-PLS-Dist), equipped with one of the proposed distance metrics as the most efficient and robust solver among the considered ones. We apply this algorithm to obtain a diverse set of high-performing solutions and analyze them to make general conclusions of the better combinations of transmission profiles. Analysis reveals that filters with large local differences in transmission improve the device performance. Moreover, the obtained top-performing solutions show significant improvement compared to the baseline.
Keywords Photonic crystal $\cdot$ Trace Gas Measurements $\cdot$ Spectrometer $\cdot$ Metaheuristic $\cdot$ Evolutionary Computation $\cdot$ Domain-Aware Optimization $\cdot$ Stochastic Combinatorial Optimization $\cdot$ Simulation Optimization
@@ 1 Introduction
As global climate change severely impacts our world, there is an increasing demand to monitor trace gases with a high spatial resolution and accuracy. At the same time, these instruments and satellites need to be compact in order to have constellations for short revisit times [1]. In this study we explore a new spectrometer instrument concept to detect trace gases where photonic crystal filters directly placed on a detector replace traditional diffraction-based optical elements [2].
@@
Trace gasses in the atmosphere slightly change the spectrum of the transmitted light due to absorption at specific wavelengths. The instrument is able to perform trace gas retrieval by measuring the total amount of reflected light each filter transmits for a certain ground pixel. By analyzing the measured intensities, it is possible determine the concentration of trace gases. The performance of the instrument, measured as accuracy and precision in gas retrieval, depends on the signal-to-noise ratio and significantly on the selected filters. Therefore, a key challenge is to select the filters that will most accurately measure the trace gas concentration. To achieve this, we simulated how 2D photonic crystals respond to light, creating a library of about 5000 different filters in various shapes and sizes. Given that the considered instrument can have at maximum 640 filters, we reduce the problem to the selection of an optimal combination of filters from the created library. In this study, we are specifically focusing on detecting methane as a test case. Therefore, we can measure the instrument’s performance by comparing the detected methane concentration with the actual concentration. While we concentrate in this work on the optimization of a trace gas imager for earth observation, the optimization algorithm introduced in this paper could be transferred to other instruments with nanostructured filter arrays, such as in biological imaging [3], mid-infrared gas detection [4] and flow cytometry [5].
@@
This paper addresses the described problem of selecting from a large library of approximately $L\approx 5\cdot 10^{3}$ filters a combination that optimizes the instrument’s performance in methane retrieval. Unlike a set, where each element is unique, filters in this selection can be chosen multiple times, i.e., a solution candidate is a multiset with $N$ elements chosen from $L$. For the remainder of this paper, we refer to this core problem as Optimal Filter Selection (OFS). The problem is addressed by an optimization algorithm that tackles both challenges: the prohibitively large number of possible solutions and the presence of noise introduced by a realistic atmospheric model. The latter is a special challenge, as many evaluations of different filter sets can result in outliers, eroding the convergence of the optimization algorithm. Therefore, it is most natural to formulate OFS as a stochastic combinatorial optimization problem.
@@ 1.1 Combinatorial Optimization
The field of combinatorial optimization [6, 7] focuses on designing algorithms to find one or more objects from a finite set that minimize a specified cost criterion.
According to a high-level classification of combinatorial optimization algorithms [7], these methods can be broadly divided into the following three categories:
  1. 1.
    Exact algorithms, which aim to compute the optimal solution with mathematical guarantees. These include classical methods such as branch-and-bound, branch-and-cut, and dynamic programming, and are extensively studied in [6, 7]. While they provide optimality, their computational cost often limits their applicability to small or moderately sized instances.
  2. 2.
    Approximation algorithms, which provide provably near-optimal solutions within a guaranteed error bound. These algorithms are especially useful for NP-hard problems where exact computation is infeasible, and are covered in depth in [8].
  3. 3.
    Heuristic and metaheuristic algorithms, including evolutionary algorithms, simulated annealing, and ant colony optimization. These are state-of-the-art numerical solvers that are widely applied in practice due to their ability to find high-quality solutions for large-scale and complex problems, albeit without formal optimality guarantees. Comprehensive reviews can be found in [9, 10].
Algorithms in the former two categories are conventional combinatorial optimization methods (CCOMs). They search via mathematical or dynamic programming and employ heuristic ideas to disregard intuitively sub-optimal solutions. Particular instances of CCOMs are implemented to solve a specific known COP and therefore heavily depend on the known structure of the addressed problem.
The third category of algorithms includes methods that apply nature-inspired meta-heuristics [11] and machine learning [12] approaches to find solution candidates with sufficiently small costs. While these optimizers do not guarantee to locate the optimal solution, they are less dependent on the constraint nature of the problem and so can be applied to different classes of COPs without major modifications. Successful applications of these methods are demonstrated in works such as [13, 14, 15]. Those and other relevant research directions are systematically studied in this recent survey [16].
@@ 1.2 Stochastic Combinatorial Optimization
Stochastic combinatorial optimization problems (SCOPs) introduce an additional layer of complexity: the cost associated with a specific solution candidate exhibits randomness, meaning it may vary upon reevaluation. This property is especially relevant for engineering applications as the cost is estimated by a simulation of a real-world process, for example, see Part IV in [17]. Application of an algorithm to a SCOP that relies on a simulation is called simulation optimization (SO) in the literature [17, 18, 19], since a closed-form mathematical expression is not available to describe the optimized function. SO can be classified as a particular case of a more general Black-Box Optimization task [20]. CCOMs can be applied as simulation optimizers, for example, see Chapter 2 of [18]. However, rapidly developing areas of relevant meta-heuristic and machine learning methods [21] contain alternative promising algorithms. As shown in Table 1 of [21], those algorithms were advantageous in a number of practical use cases. See the survey [22] for more examples and a fresh systematic review of the related publications.
@@ 1.3 Addressing OFS problem
While CCOMs for OFS might exist, the complexity of deriving them may not be practical. Therefore, we leverage the efficiency of numerical methods, such as meta-heuristics and machine learning, to obtain a high-quality approximate solution to OFS. Meta-heuristics require neither explicit analytical models nor gradients, and have a long track record of delivering high-quality solutions on complex, noisy, and nonconvex problems across many application domains as discussed in Sec. 1.11.2. They are also flexible and easily extensible because they are robust to changes in the simulator or constraint set. Therefore, addressing the OFS problem with meta-heuristics is the best available approach, which we employ in this work.
Given the vast number of existing general-purpose methods, we use the latest (known to us) comprehensive survey [19]. This survey provides a clear description of various methods applicable to SO, helping us in selecting optimizers for OFS task. Independent of our choice, it is guaranteed by the No Free Lunch (NFL) theorem that none of the selected optimizers is the best on average for a big enough class of problems, meaning, a class of problems closed under the permutations of the search space [23]. Therefore, knowing the best optimizer for OFS in advance is intractable. Hence, we follow the approach suggested in Section 2.7 of [10]. We conduct experiments where we repeatedly approximate the solution of the OFS problem using every chosen algorithm. This allows us to statistically rank the performance of each algorithm and identify the most suitable ones for our specific problem. We call the top-performing solvers the leading algorithms.
@@
    Compare meta-heuristics that represent various classes of SCOP solvers, and select the solvers with the best performance on OFS;
@@
    To implement a key step required for obtaining an effective mutation operator in this distance-driven evolutionary framework, we design a heuristic method, called DDA-EA, to approximately solve the inverse of the Linear Assignment Problem;
@@
    We introduce a novel method, UMDA-U-PLS-Dist, which incorporates a user-defined distance metric to the Estimation of Distribution Algorithm to guide the search process more effectively, enhancing the exploitation of problem-specific structure.
  5. 5.
    We obtain a diverse set of high-performing solutions to the OFS problem and analyze them to make general conclusions of the better combinations of transmission profiles.
@@ 2 Description of the Trace Gas Measurement Device
[…]
@@
As the light is incident on the filters, different integrated intensities for each filter are measured for a single ground resolution element. From this detector data, the gas concentrations can be retrieved. The retrieval model is described in previous work [24]. In short, the trace concentration is used as input for the atmosphere model, see Figures 1 and 2. The atmosphere model produces the spectrum reflected by the Earth, which is used as input for the instrument model, taking into account the specific filters, which compute the measured values to which the photon and read noise are added. From these noisy measurement values, an inverse model of the instrument and atmosphere are used to retrieve the trace gas concentration using a fitting routine. By performing this fit for different noisy realizations, the fit precision of the trace gas concentration is assessed. This precision is the performance indicator of the instrument concept and should be as low as possible.
In general, the achieved gas concentration retrieval precision depends on many factors such as optical aperture and ground resolution. The specifications and parameters defining the final performance of the instrument are chosen to be reasonable and achievable. Therefore, the achieved precision after filter selection optimization is a representative and likely achievable retrieval precision. Currently, the expected retrieval error of TANGO (Twin Anthropogenic Greenhouse gas Observers), an ESA Scout mission to measure methane, which consists of a traditional state-of-the-art methane retrieval instrument, is below 1.0% [25]. However, the optical design of the proposed instrument using the photonic crystals is expected to be 5-10 times smaller.
Prior to the development of optimization algorithms, we constructed a reasonable filter selection, which we will refer to as SRON baseline. As this initial solution, we choose filters based on the sharpness of the spectral transmission features. We used the second moment of the Fourier transform of the transmission profile to rank all filters. Filters with sharp spectral features would result in large amplitudes for the high-frequency components. From this ranked list, 160 filters are chosen with the largest second moment, and this filter set is used as a reference solution.
@@ 3.1 Search Space
The search space is defined as sequences of size $N=640$ of the filters from the library $\mathbb{L},L\coloneqq|\mathbb{L}|=4374.$ One filter can be included multiple times in the sequence. Let us denote the space of all such sequences as $\mathbb{L}^{N}$. We assume that the cost criterion is given as a function $F_{N}:\mathbb{L}^{N}\to\mathbb{R}.$ Our goal is to minimize this criterion, meaning find such a sequence $\bm{x}^{*}\in\mathbb{L}^{N}$ that $F_{N}\!\left(\bm{x}^{*}\right)=\min\!\left\{F_{N}(\bm{x})\,\middle|\,\bm{x}\in\mathbb{L}^{N}\right\}.$
Although heuristic algorithms are specifically designed to navigate large and complex search spaces without exhaustive enumeration, their effectiveness is still fundamentally limited by the size and structure of the search space. In extremely large combinatorial spaces such as $\mathbb{L}^{N}$, even heuristics face difficulties due to several factors: the high likelihood of getting trapped in local optima, the sparsity of high-quality solutions relative to the total number of candidates, and the limited number of evaluations typically allowed due to computational constraints. As a result, the probability of encountering a solution of reasonably good quality becomes increasingly small as $N$ grows, especially in the absence of strong guidance mechanisms. This motivates our restriction to a smaller subspace $\mathbb{L}^{M}$, with $M\ll N$, where heuristic search becomes more tractable while still retaining sufficient diversity for effective optimization. Limiting the number of filter types provides a practical benefit: reduced manufacturing costs due to a less complex production process for a smaller variety of different filters, as discussed in Section 2.
We encode filters in the set $\mathbb{L}$ by integers $1,2,\ldots,L.$ Given a vector of $M$ filters, we create a vector of $N$ filters by repeating each filter multiple times. More precisely, if $k\coloneqq\lfloor N/M\rfloor$ and $r\coloneqq N-k\cdot M,$ then the first $r$ filters are repeated $k+1$ time and the rest MrM-r filters are repeated $k$ times. We can see that the size of the final selection is exactly $N$, indeed $r\cdot(k+1)+(M-r)\cdot k=r+k\cdot M=N-k\cdot M+k\cdot M=N.$ For example, if $N=5,M=3$, then $k=1,r=2.$ If the selection of size $M$ is $(2,4,7)$, then the restored selection of size $N$ is $(2,2,4,4,7).$ Note that $r=2$ first filters, meaning filters $2$ and $4$, are repeated $k+1=2$ times and the last filter $7$ is repeated $k=1$ time.
Let us denote the mapping that does this transformation as $\Phi:\mathbb{L}^{M}\to\mathbb{L}^{N}.$ Given a vector of filters $\bm{x}\in\mathbb{L}^{M}$ we can assess its cost by $\mathrm{F}:\mathbb{L}^{M}\to\mathbb{R},\mathrm{F}\coloneqq\mathrm{F}_{N}\circ\Phi.$ Notation $f\circ g(x)$ stands for $f(g(x))$. Now, we are ready to define the problem in the form, which is addressed in this paper.
Problem Formulation 1 (OFS problem in the form addressed in this work).
@@
For the given constant $M$, let $\mathscr{L}=\left\{\bm{x}=(x_{1},x_{2},\dots,x_{M})\,\middle|\,x_{i}\in\mathbb{L}\right\}$ be the space of candidate solutions. For function $\mathrm{F}:\mathscr{L}\to\mathbb{R}$ such that $\mathrm{F}(\bm{x})=\mathrm{F}_{N}\circ\Phi(\bm{x})$, OFS problem is to find $\bm{x}^{*}\in\mathscr{L}$ which satisfies the Eq. (1):
@@
The cardinality of the space $\mathscr{L}$ equals $L^{M}$, which is large in our particular instance of OFS: $L\approx 5000,M=16,L^{M}\approx 10^{59}$. However, all permutations of components in a candidate solution $\bm{x}\in\mathscr{L}$ result in the same vector, so they all have the same associated cost. We show how to exploit this property for numerical optimization in Section 5.
The OFS in the form defined in Formulation 1 is convenient for heuristic optimization, since every component in a candidate solution can be considered independently from others. From a practical perspective, the constraint in Eq. (1b) is soft, meaning that minor violations are acceptable. This is because solutions with an arbitrary number of filters can still be manufactured in practice. The restriction to exactly $M$ filters is introduced primarily to reduce the combinatorial complexity of the search space, rather than to enforce a strict physical limitation.